# Mesh-clustered Gaussian process emulator for partial differential equation boundary value problems

### 02 Feb 2024 v4

@@ 2 PDE boundary value problems and mesh-based numerical methods
where $\Omega$ is a compact domain in $\mathbb{R}^{d}$ with a Lipschitz boundary denoted by $\partial\Omega$, $\mathcal{L}(\cdot;\mathbf{x})$ and $\mathcal{G}(\cdot;\mathbf{x})$ are differential operators on $\Omega$ and $\partial\Omega$ given the input $\mathbf{x}$, respectively, $f(\cdot;\mathbf{x})$ and $g(\cdot;\mathbf{x})$ are two known functions on $\Omega$ and $\partial\Omega$ with the input $\mathbf{x}$, and $u(\mathbf{s})$ is the solution to the partial differential equations (1). Given the input $\mathbf{x}$, the exact solution to the equations when uniqueness holds, denoted by $u_{0}(\mathbf{s};\mathbf{x})$, usually cannot be written explicitly. Instead, numerical methods are used to approximate the exact solution, for which mesh-based numerical methods are most widely used, including the finite difference method (FDM), finite volume method (FVM), and finite element method (FEM) (Brenner and Scott,, 2007; Tekkaya and Soyarslan,, 2019). The extensions to mesh-free methods, such as the collocation method (Golberg et al.,, 1999; Wendland,, 1998; Fasshauer,, 1999; Wendland,, 1999; Fasshauer,, 1996), are straightforward and will be discussed in the remarks. Specifically, mesh-based numerical methods subdivide a large system into smaller, simpler parts by a particular space discretization in the space dimensions, which is implemented by the construction of a mesh of the object. Figure 1 demonstrates a mesh specification for a 2-dimensional FEM problem, in which the triangular elements (Figure 1C) connect all characteristic points (called nodes) that lie on their circumference, and these connections are mathematically expressed through a set of functions called shape functions.
Figure 1: Introduction of finite element method (modified from Tekkaya and Soyarslan, (2019)).
@@
where $v_{j}(\mathbf{s})$ is the (given) shape function depending on the discretization, and $\beta_{j}(\mathbf{x})$ is the corresponding coefficient, which is independent of $\mathbf{s}$. The shape function has the Kronecker Delta property, that is, $v_{j}(\mathbf{s}_{i})=1$ if $i=j$ and $v_{j}(\mathbf{s}_{i})=0$ if $i\neq j$. This ensures that the function interpolates the solution at the mesh nodes, i.e., $u_{N}(\mathbf{s}_{j};\mathbf{x})=\beta_{j}(\mathbf{x})$ for $j=1,\ldots,N$. In other words, the solution of the PDE at the mesh node $\mathbf{s}_{j}$ is equal to $\beta_{j}(\mathbf{x})$. As an example, for a triangular element with six nodes as in Figure 1C, the quadratic shape functions, which will be adopted in our later implementation in Section 5, can be defined as:

### 14 Feb 2024 v5

@@ 2 PDE boundary value problems and mesh-based numerical methods
where $\Omega$ is a compact domain in $\mathbb{R}^{d}$ with a Lipschitz boundary denoted by $\partial\Omega$, $\mathcal{L}(\cdot;\mathbf{x})$ and $\mathcal{G}(\cdot;\mathbf{x})$ are differential operators on $\Omega$ and $\partial\Omega$ given the input $\mathbf{x}$, respectively, $f(\cdot;\mathbf{x})$ and $g(\cdot;\mathbf{x})$ are two known functions on $\Omega$ and $\partial\Omega$ with the input $\mathbf{x}$, and $u(\mathbf{s})$ is the solution to the partial differential equations (1). Given the input $\mathbf{x}$, the exact solution to the equations when uniqueness holds, denoted by $u_{0}(\mathbf{s};\mathbf{x})$, usually cannot be written explicitly. Instead, numerical methods are used to approximate the exact solution, for which mesh-based numerical methods are most widely used, including the finite difference method (FDM), finite volume method (FVM), and finite element method (FEM) (Brenner and Scott,, 2007; Tekkaya and Soyarslan,, 2019). The extensions to mesh-free methods, such as the collocation method (Golberg et al.,, 1999; Wendland,, 1998; Fasshauer,, 1999; Wendland,, 1999; Fasshauer,, 1996), are straightforward and will be discussed in the remarks. Specifically, mesh-based numerical methods subdivide a large system into smaller, simpler parts by a particular space discretization in the space dimensions, which is implemented by the construction of a mesh of the object. Figure 1 demonstrates a mesh specification for a 2-dimensional FEM problem, in which the triangular elements connect all characteristic points (called nodes) that lie on their circumference, and these connections are mathematically expressed through a set of functions called shape functions.
Figure 1: Introduction of finite element method; similar to one from Tekkaya and Soyarslan, (2019).
@@
where $v_{j}(\mathbf{s})$ is the (given) shape function depending on the discretization, and $\beta_{j}(\mathbf{x})$ is the corresponding coefficient, which is independent of $\mathbf{s}$. The shape function has the Kronecker Delta property, that is, $v_{j}(\mathbf{s}_{i})=1$ if $i=j$ and $v_{j}(\mathbf{s}_{i})=0$ if $i\neq j$. This ensures that the function interpolates the solution at the mesh nodes, i.e., $u_{N}(\mathbf{s}_{j};\mathbf{x})=\beta_{j}(\mathbf{x})$ for $j=1,\ldots,N$. In other words, the solution of the PDE at the mesh node $\mathbf{s}_{j}$ is equal to $\beta_{j}(\mathbf{x})$. As an example, for a triangular element with six nodes as in Figure 1, the quadratic shape functions, which will be adopted in our later implementation in Section 5, can be defined as: